# Homogeneous maximizers of the Blaschke–Santaló-type functionals

### v1

@@ Abstract.
We study Blaschke–Santaló-type inequalities for $N\geq 2$ sets (functions) and a special class of cost functions. In particular, we prove new results about reduction of the maximization problem for the Blaschke–Santaló-type functional to homogeneous case (functional inequalities on the sphere) and extend the symmetrization argument to the case of $N>2$ sets. We also discuss links to the multimagrinal optimal transportation problem and the related sharp transportation-information inequalities.
@@ 1. Introduction
We prove the following generalization of the both versions of the Blaschke–Santaló inequality.
@@
Let $N\geq 2$ be a natural number. Consider the following cost function
@@
In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
This example can be considered as an illustration of several results of this work, where the central question is the relation between the functional and the set version of the Blaschke–Santaló type functional. We can put it into the following form: when the Blaschke–Santaló functional admits homogeneous extremizers for given homogeneous reference measures and homogeneous cost function?
@@
    There are several extensions of the Blaschke–Santaló inequality for special types of the generalized Legendre transform. See, for instance, Rotem [35], Artstein–Slomka [3], Artstein–Sadowski–Wyczesany [2], Fradelizi–Gozlan–Sadovsky–Zugmeyer [18].
@@
    Kolesnikov and Werner [27] proved that (6) holds under assumption of unconditionality. Some generalizations of this result were obtained by Kalantzopoulos and Saroglou [23]. In particular, they conjectured the following set version of (6):
@@
Another central result of this work is the following Theorem about homogeneity of maximizers of the Blaschke–Santaló functional.
@@
In section 2 we discuss the relation of the Blaschke–Santaló inequality to the optimal transportation problem. We show in sections 4-5 that $(p_{1},\cdots,p_{N})$-homogeneous cost functions admit other nice properties. In particular, we prove that the corresponding functional and set versions of the Blaschke–Santaló inequality for these cost functions are equivalent. In addition, we show that the Blaschke–Santaló functional increases under the appropriate version of the Steiner symmetrization procedure.
@@ Acknowledgements
The author was supported by RSF project 25-11-00007. The author thanks Matthieu Fradelizi for his interest and comments about the bibliography.
@@ 3. Homogeneity of the maximizers
Let $c\geq 0$ be sufficiently regular cost function on $(\mathbb{S}^{n-1})^{N}$, $\varphi_{i}$ be nonnegative functions satisfying (27) and, in addition, $\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}<\infty$ for all $1\leq i\leq N$. Consider mass transportation problem with cost function $\log c$ and probability measures
@@ 4. Inequalities for sets and for functions
In this section we continue our study of homogeneous solutions and find a direct link between the functional and the set inequality. Let us specify assumptions on $c$ in this section.
@@ 5. Symmetrization
We will assume in this section that the polar $c$-transform preserves symmetric and convex sets, i.e.
$B_{j}=\{x_{j}:c(x_{1},\cdots,x_{N})\leq 1,\forall x_{i}\in B_{i},i\neq j\}$
@@
    $(y_{i_{1}},t_{i_{2}})\to c((y_{1},t_{1}),\cdots,(y_{N},t_{N}))$
    is convex for all fixed $y_{i_{k}}$, $k\neq 1$ and $t_{i_{j}}$, $j\neq 2$
@@
$1$ $\geq\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
$=\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},-s),(b,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
$\geq c\bigl((y_{1},\frac{t-s}{2}),(\frac{a+b}{2},r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr).$
@@ 6. Main example
for some Borel function $\rho$. Then
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(x_{i})dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\bigl(\sum_{j=1}^{n}t^{A}_{j}\bigr)dt\Bigr)^{A},$
@@
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}e^{-\frac{1}{A}\sum_{j=1}^{n}t^{A}_{j}}dt\Bigr)^{A}.$
@@
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})dx_{i,1}\cdots dx_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(e^{t_{i}})e^{\sum_{j=1}^{n}t_{i,j}}dt_{i,1}\cdots dt_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\prod_{i=1}^{N}f_{i}^{{\frac{1}{A\alpha_{i}}}}(e^{t_{i}})e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{\sum_{i=1}^{N}\frac{t_{i,j}}{\alpha_{j}}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
$=\Bigl(\int_{\mathbb{R}^{n}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{As_{j}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}=\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}t^{A}_{j}\Bigr)dt\Bigr)^{A}.$
Remark 6.2.
The statement remains true if we take $(x_{i,1}\cdots x_{i,n})^{\gamma}dx_{i}$, $\gamma>-1,$ for the reference measure instead of the Lebesgue measure.
We prove the main result of this paper.
@@
(Theorem 1.1) Let $N\geq 2$ be a natural number. Consider the following cost function
@@
$\prod_{i=1}^{N}|A_{i}|\leq|B_{N}|^{N}.$
In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
@@
Clearly, $c$-polar transform preserves symmetric and unconditional sets and according to Example 5.2 the cost function satisfies $(j,i_{1},i_{2})$-assumption for all indices. By Corollary 5.4 the desired inequality can be reduced the unconditional case. Thus without loss of generality one can assume that $A_{i}$ are unconditional. Then the set inequality follows from Proposition 6.1 (we take $f_{i}=I_{A_{i}}$ and $\rho(t)=I_{[0,1]}(t)$). The functional inequality follows from Theorem 4.4. The statement about homogeneity of maximizers follows from Corollary 3.4. Finally, description of maximizers in terms of Minkowski functionals follows from Corollary 4.5. ∎
@@
where $\alpha_{i}\leq 1$. Let $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ be $N$ symmetric convex sets satisfying $c(x_{1},\cdots,x_{N})\leq 1$, provided $x_{i}\in A_{i}$, $1\leq i\leq N$. Then
$\prod_{i=1}^{N}|A_{i}|^{\frac{1}{\alpha_{i}}}\leq|B|^{A},$
@@
Let $\{V_{i}\}$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$. Then
$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{A}\sum_{i=1}^{n}|y_{i}|^{A}}dy\Bigr)^{A}.$
In addition, every maximizing tuple $\{\Phi_{i}\}$ with $\Phi_{i}(0)=0$ consists of $A$-homogeneous functions of the form $\{\frac{\frac{1}{\alpha_{i}}}{A}\|x_{i}\|^{A}_{K_{i}}\}$, where $\{K_{i}\}$ is a maximizing tuple of the corresponding set problem.
7. Concluding remarks and open questions
In view of Theorems 1.1, 6.4 it is natural to ask about examples of the Blaschke–Santaló inequality for non-Lebesgue reference measures. One such example can be constructed from the result of Theorem 6.4 just by change of variables
@@ 7. Concluding remarks and open questions
where $p>1,\frac{1}{p}+\frac{1}{q}=1$ and $\rho$ is $\frac{2-p}{p-1}$-homogeneous can be found in [11]. Another remarkable example is given in Fradelizi–Meyer–Zvavitch [20] (Theorem 7):
(32) $\mu(A)\mu(A^{\circ})\leq\mu(B_{2})^{2}$
for any convex symmetric $A$ and unconditional log-concave measure $\mu$. We remind the reader, that Cordero-Erausquin conjectured in [12] that (32) is true for any symmetric log-concave $\mu$. It is possible to extend the result of [20] to the case of $N>2$ sets and functionals of the type $\prod_{i=1}^{N}\mu(A_{i})$ and the cost function $c=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$ using symmetrization and an adapted variant of Proposition 6.1, but we don’t pursue this here. Unlike this, it is problematic to extend inequality (31) to the case of $N>2$ sets using symmetrization. This is because we need log-concavity of measures for symmetrization and this assumption does not fit well with homogeneity.
Finally, we observe that the barycentric cost function $c_{bar}(x)=\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ does not satisfy the key assumption of $(p_{1},\cdots,p_{N})$-homogeneity and our results can not be applied to this case. We know by the result of Nakamura and Tsuji [32] that the corresponding functional $\prod_{i=1}^{N}\int e^{-V_{i}}dx_{i}$ with constraints $\sum_{i=1}^{N}V_{i}(x_{i})\geq\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ admits $2$-homogeneous (even quadratic) extremizers, but nothing is known about solutions to the corresponding set problem and about the relations between both problems (uniqueness, homogeneity of any maximizing tuple etc).

### v2

@@ Abstract.
We study Blaschke–Santaló-type inequalities for $N\geq 2$ sets (functions) and a special class of cost functions. In particular, we prove new results about reduction of the maximization problem for the Blaschke–Santaló-type functional to the homogeneous case (functional inequalities on the sphere) and prove new inequalities for $N>2$ sets by symmetrization argument. We also discuss links to the multimagrinal optimal transportation problem and the related sharp transportation-information inequalities.
@@ 1. Introduction
The following interesting example of the Blaschke–Santaló inequality was proved by Kalantzopoulos and Saroglou in [23]
@@
(Kalantzopoulos, Saroglou, [23]) Let $N\geq 2$ be a natural number. Consider the following cost function
@@
We extend this result in section 6 of our work, where we get, in particular a generalization for functionals of the type $\prod_{i=1}^{N}|A_{i}|^{\frac{1}{\alpha_{i}}}$ . This example can be considered as an illustration of several results of this work, where the central question is the relation between the functional and the set version of the Blaschke–Santaló type functional. We can put it into the following form: when the Blaschke–Santaló functional admits homogeneous extremizers for given homogeneous reference measures and homogeneous cost function?
@@
    There are several extensions of the Blaschke–Santaló inequality for special types of the generalized Legendre transform. See, for instance, Rotem [35], Artstein–Slomka [3], Artstein–Sadovski–Wyczesany [2], Fradelizi–Gozlan–Sadovsky–Zugmeyer [18].
@@
    Kolesnikov and Werner [27] proved that (6) holds under assumption of unconditionality. Some generalizations of this result were obtained by Kalantzopoulos and Saroglou [23]. In particular, the following version of (6) was conjectured:
@@
The central result of this work is the following Theorem about homogeneity of maximizers of the Blaschke–Santaló functional.
@@
In section 2 we discuss the relation of the Blaschke–Santaló inequality to the optimal transportation problem. Theorem 1.3 is proved in section 3. We show in sections 4-5 that $(p_{1},\cdots,p_{N})$-homogeneous cost functions admit other nice properties. In particular, we prove that the corresponding functional and set versions of the Blaschke–Santaló inequality for these cost functions are equivalent. In addition, we discuss the monotonicity of the Blaschke–Santaló functional under the appropriate version of the Steiner symmetrization procedure. In section 6 we consider examples.
@@ Acknowledgements
The author was supported by RSF project 25-11-00007. The author thanks Matthieu Fradelizi for his interest and comments about the bibliography and Christos Saraglou for comments about Theorem 1.1.
@@ 3. Homogeneity of the maximizers
Let $c\geq 0$ be sufficiently regular cost function on $(\mathbb{S}^{n-1})^{N}$, $\varphi_{i}$ be nonnegative functions satisfying (27) and, in addition, $\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}<\infty$ for all $1\leq i\leq N$. Consider the mass transportation problem with the cost function $\log c$ and probability measures
@@ 4. Inequalities for sets and for functions
In this section we continue our study of homogeneous solutions and find a direct link between the functional and the set inequality. This was partially done in the previous sections, where we find conditions ensuring homogeneity of the maximizers in the functional inequality. Here we discuss how to recover the functional inequality from the set inequality using ”lawer-cake” representation.
Assumption on reference measures:
$m_{i}=\rho_{i}(x_{i})dx_{i},$
where $\rho_{i}$ is $r_{i}$-homogeneous.
@@ 5. Symmetrization
Remark 5.1.
Unlike sections 3-4 we don’t assume homogeneity of the cost function $c$ and the reference measures.
Our approach closely follows the arguments of Kalantzopoulos and Saraglou from [23], in particular, our important $(j,i_{1},i_{2})$-assumption below is satisfied by the cost functions considered in [23]. However we establish our result in a more general situation. In particular, we consider more general class of functions and the reference measures.
@@
  1. (1)
    The polar $c$-transform preserves symmetric and convex sets, i.e.
    $B_{j}=\{x_{j}:c(x_{1},\cdots,x_{N})\leq 1,\forall x_{i}\in B_{i},i\neq j\}$
  2. (2)
@@
    $c_{i_{1},i_{2}}(y_{i_{1}},t_{i_{2}})=c((y_{1},t_{1}),\cdots,(y_{N},t_{N}))$
    has convex level sets for all fixed $y_{i_{k}}$, $k\neq 1$ and $t_{i_{j}}$, $j\neq 2$
  3. (3)
@@
$c\bigl((y_{1},-s),(b,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})=c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\leq 1$
by item (3) of $(n,1,2)$ assumption (see Definition 5.2), hence by item (2)
$1\geq c\bigl((y_{1},\frac{t-s}{2}),(\frac{a+b}{2},r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr).$
@@ 6. Main example
In this section we discuss examples of the Blaschke–Santaló-type inequalities, generalizing Theorem 1.1 from [23]. Our extensions go in the following directions: 1) extension to unconditional log-concave reference measures in the spirit of Theorem 7 from [20], 2) extension to product functionals with different weights, i.e. functionals of the type $\prod_{i=1}^{N}m^{\frac{1}{\alpha_{i}}}(A_{i})$.
For the case $N=2$ the following result was obtained in Fradelizi–Meyer–Zvavitch [20] (Theorem 7):
(31) $\mu(A)\mu(A^{\circ})\leq\mu(B_{2})^{2}$
for any convex symmetric $A$ and unconditional log-concave measure $\mu$. We remind the reader, that Cordero-Erausquin conjectured in [12] that (31) is true for any symmetric log-concave $\mu$.
@@
for some Borel function $\rho$. Let $m=e^{-V}dx$ be a measure on $\mathbb{R}^{n}_{+}$ such that the function $(t_{1},\cdots,t_{n})\to V(e^{t_{1}},\cdots,e^{t_{n}})$ is convex.
Then
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(x_{i})m(dx_{i})\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\bigl(\sum_{j=1}^{n}t^{A}_{j}\bigr)m(dt)\Bigr)^{A},$
@@
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}e^{-\alpha_{i}V_{i}(x_{i})}m(dx_{i})\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}e^{-\frac{1}{A}\sum_{j=1}^{n}t^{A}_{j}}m(dt)\Bigr)^{A}.$
@@
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})e^{-V(x_{i})}dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}$ $=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})e^{-V(x_{i})}dx_{i,1}\cdots dx_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
$=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(e^{t_{i}})e^{-V(e^{t_{i}})}e^{\sum_{j=1}^{n}t_{i,j}}dt_{i,1}\cdots dt_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\prod_{i=1}^{N}f_{i}^{{\frac{1}{A\alpha_{i}}}}(e^{t_{i}})e^{-\frac{1}{A\alpha_{i}}V(e^{t_{i}})}e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{\sum_{i=1}^{N}\frac{t_{i,j}}{\alpha_{j}}}\Bigr)e^{-V(e^{\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i}}{\alpha_{i}}})}e^{\sum_{k=1}^{n}s_{k}}ds\Bigr)^{A}$
$=\Bigl(\int_{\mathbb{R}^{n}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{As_{j}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}e^{-V(e^{s})}ds\Bigr)^{A}=\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}t^{A}_{j}\Bigr)e^{-V(t)}dt\Bigr)^{A}.$
@@
(Extension of Theorem 1.1) Let $N\geq 2$ be a natural number and $m$ be a log-concave unconditional measure on $\mathbb{R}^{n}$. Consider the following cost function
@@
$\prod_{i=1}^{N}m(A_{i})\leq m(B_{N})^{N}.$
Let $\{V_{i}\}$, $1\leq i\leq N,$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(m)$ for every $i$. Then
$\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}m(dx_{i})\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}m(dy)\Bigr)^{N}.$
In addition, let $m$ be the Lebesgue measure and $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
@@
Clearly, $c$-polar transform preserves symmetric and unconditional sets and according to Example 5.3 the cost function satisfies $(j,i_{1},i_{2})$-assumption for all indices. By Corollary 5.5 the desired inequality can be reduced to the unconditional case. Thus without loss of generality one can assume that $A_{i}$ are unconditional. The same symmetrization procedure can be done for functions (we don’t present the complete proof here, but we can refer to [11] Theorem 4.1, where the symmetrization for functions was considered in detail for $N=2$. The case $N>2$ follows the arguments of Lemma 5.4 and Theorem 4.1 of [11]). Finally, after reduction to the unconditional case, both inequalities (for sets and functions) follow from Proposition 6.1. The statement about homogeneity of maximizers follows from Corollary 3.4. Finally, description of maximizers in terms of Minkowski functionals follows from Corollary 4.5. ∎
@@
where $\alpha_{i}\leq 1$. Let $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ be $N$ symmetric convex sets satisfying $c(x_{1},\cdots,x_{N})\leq 1$, provided $x_{i}\in A_{i}$, $1\leq i\leq N$ and $m$ be a log-concave unconditional measure. Then
$\prod_{i=1}^{N}m(A_{i})^{\frac{1}{\alpha_{i}}}\leq m(B)^{A},$
@@
Let $\{V_{i}\}$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(m)$ for every $i$. Then
$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}m(dx_{i})\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{A}\sum_{i=1}^{n}|y_{i}|^{A}}m(dy)\Bigr)^{A}.$
In addition, let $m$ be the Lebesgue measure. Then every maximizing tuple $\{\Phi_{i}\}$ with $\Phi_{i}(0)=0$ consists of $A$-homogeneous functions of the form $\{{\frac{1}{A\alpha_{i}}}\|x_{i}\|^{A}_{K_{i}}\}$, where $\{K_{i}\}$ is a maximizing tuple of the corresponding set problem.
Remark 6.4.
In view of Theorems 1.1, 6.3 it is natural to ask about examples of the Blaschke–Santaló inequality for homogeneous non-Lebesgue reference measures. One such example can be constructed from the result of Theorem 6.3 with Lebesgue reference measure just by change of variables
@@
where $p>1,\frac{1}{p}+\frac{1}{q}=1$ and $\rho$ is $\frac{2-p}{p-1}$-homogeneous can be found in [11]. We observe that it is problematic to extend inequality (32) to the case of $N>2$ sets using symmetrization in the homogeneous case. This is because we need log-concavity of measures for symmetrization and this assumption does not fit well with homogeneity.